Two cars A and B are traveling toward each other at

speeds of 45 km/hr and 70 km/hr respectively. If A

is blowing its horn, what is the relative pitch heard

by a passenger in B, considering that the velocity of

sound is 344 m/s.



1
Expert's answer
2021-04-01T18:41:57-0400
"\frac{f}{f_0}=\frac{c+v_r}{c-v_s}\\\frac{f}{f_0}=\frac{344+\frac{70}{3.6}}{344-\frac{45}{3.6}}
\\\frac{f}{f_0}=1.096"